### v1

@@ Abstract
The details of multiple charge transfer within or among molecules (including the accompanying molecular structure evolution and energy distribution) are typically not accessible on a single molecule level in experiments targeting complex condensed chemical or biological systems. In order to gather such detailed insight, small prototype systems that cover the essence of such processes need to be identified and investigated. Here, we employ a small system consisting of a combination of covalent and van der Waals bonds for our studies, namely N2Ar dimers. We use synchrotron radiation to site-selectively enable the charge transfer processes and perform a coincidence measurement of the resulting electrons and ions. In combination with ab initio calculations, this approach enables a step-by-step tracking of the charge transfer and fragmentation dynamics. We find that ultrafast structural evolution of the dimer can trigger a second CT, thereby opening complex reaction pathways, in which electrons transfer back and forth between Ar and N2, and nonadiabatic transitions occur twice through conical intersections. These results demonstrate that multiple CT-induced transitions, particularly in such a simple dimer system, provide benchmark insights into the mechanisms of nonadiabatic reactions in complex natural systems.
1 Introduction
@@
Ultrafast charge transfer (CT), which redistributes electronic charge within or between molecular units, is one of the fundamental processes that underlie the emergence of function in natural and engineered systems [1, 2, 3, 4, 5, 6, 7]. Charge and energy transport span length scales from small dimers, simple and complex molecules in the gas or liquid phase to photoactive materials and proteins [8, 9, 10, 11, 12, 13]. Tracing complex reactions induced by CTs is usually inaccessible via bulk ensemble measurements and the details of such vital processes remain poorly understood. Direct observation of (multiple) CTs at the single-molecule level is therefore essential, as it can offer unique insights into the process’ transient stages and unravel the most elementary quantum dynamics of both electrons and nuclei that participate in the complex process of charge transfer[14, 15].
@@
In reality, the exchange of charge and energy occurs typically in complex systems, where covalent interactions and van der Waals bonding coexist normally. The van der Waals bond, the weakest noncovalent interaction, is ubiquitous in nature, governing the properties of nearly all organic solids and liquids [16, 17, 18]. It is responsible for phenomena such as the low boiling and melting points of substances, the layered structure of graphite, and the folding of DNA helices and proteins [19, 20], where multiple CTs across the interfaces of layers or molecules following optical excitation play a fundamental role in the regulation of optoelectronic responses. Prototype systems for van der Waals bonds consist in many cases of rare gases. Simple rare-gas dimers, however, are not an ideal candidate to study the above mentioned reactions. Multiple CTs are even unexpected in these systems, since they tend to dissociate already at the instant of the first CT [21, 22]. A suitable prototype system for our studies is the N2Ar dimer. It features a strong covalent bond between the two nitrogen atoms and a weak van der Waals bond between the argon atom and the N2 molecule. Its distinct properties make it a prototypical emulating model for experiment and quantum-chemical calculations to probe reaction mechanisms relevant to complex systems. First, the nearly matching ionization potentials [23, 24] of its two constituents make intermolecular reactions more effective under photoionization and photoexcitation scenarios by creating numerous crossings on the potential energy surfaces. Second, its structural flexibility, especially the stretching of its bonds [25] enables access to a manifold of nonadiabatic pathways through conical intersections and initiates the occurrence of multiple CT. Therefore, a system as simple as N2Ar can effectively mimic the ensemble reactions occurring in complex condensed chemical and biological systems, with the Ar atom representing species of interest and the N2 molecule the environment.
@@
X-ray-induced inner- or core-shell ionization offers an element-specific route to prepare this system in well-defined charge states. In this work, we investigate experimentally and theoretically the photoionization-induced nonadiabatic dynamics in N2Ar, focusing on ultrafast multiple CT processes. Using datasets recorded at different photon energies (two in the X-ray and one in the XUV regime), we selectively initiate double ionization (DI) of either N2 or Ar via Auger decay or direct double ionization, thereby preparing well-defined initial states. By analyzing the momenta of dimer-fragments and electrons in the final state, we have unambiguously disentangled various CT pathways, enabling reconstruction of fragmentation pathways driven by SCT and double CT (DCT). Our findings demonstrate that even in such a simple dimer multiple CT-induced transitions can drive complex reaction channels, providing fundamental insight into charge redistribution as a core nonadiabatic mechanism underlying functional processes in complex natural systems.
@@
2 Expected charge transfer channels and involved potential energy surfaces
@@
To gain intuitive insight into how the observed reaction channels – accompanied by ultrafast electron emission and CTs – are initiated and evolve from the doubly charged N2Ar dimer, we calculated the corresponding potential energy surfaces (PESs) and traced the possible fragmentation pathways. Details on our calculations are provided in the Methods section. For [N2Ar]2+, charge localization can occur in three distinct configurations: NAr, N2Ar2+, and NAr+. The high density of electronic states, together with the vibrational and rotational degrees of freedom of N2, produces significant overlap among the PESs of these configurations across a wide range of N2Ar bond lengths and angles. This extensive overlap facilitates exceptionally rapid nonadiabatic CT dynamics, a phenomenon rarely encountered in simpler systems such as Ar2 dimers. Selective ionization with X-ray photons of distinct energies enables preparation of specific initial states: removal of an N 1s or Ar 2p electron, followed by Auger decay, generates NAr states (e.g., , a, ) or N2Ar2+ states (e.g., , , ) [31, 32]. Subsequent intermolecular interactions drive structural evolution of the cluster ion, potentially leading to a nearly linear N2Ar geometry.
@@
Figure 1 presents the two-dimensional PESs for three electronic states of the linear [N2Ar]2+ ion: (N(a)Ar, NAr2+(), and N(D)Ar+()). When the system is initially prepared in either N(a)Ar or N2Ar2+(), the N-N vibrational wave packet lies within the crossing region with the N(D)Ar+() surface. This enables prompt SCT, producing the ion pair N + Ar+. Because the equilibrium bond length of N(D) is substantially larger than those of neutral N2 or N(a), the N-N bond elongates. At the same time, Coulomb repulsion in the pair N + Ar+ drives a stretching of the N2-Ar bond, thereby promoting dissociation along two distinct SCT pathways:
N2Ar2+ / NAr NAr+ N + Ar+ I
N2Ar2+ / NAr NAr+ N+ + N + Ar+ II
@@
As the N-N bond continues to elongate, the PESs of N(D)Ar+() and N(a)Ar intersect once more, enabling a second CT event. For clusters initially prepared in the N(a)Ar state, dissociation proceeds via a two-step sequence: first, N(a) captures an electron from Ar, forming NAr+; subsequently, as the N-N bond stretches, an electron is transferred back to Ar+, overcoming a substantial potential barrier. In contrast, for clusters initially in the N2Ar2+() state, neutral N2 sequentially accepts two electrons from Ar2+, ultimately driving dissociation. Remarkably, if the N2-Ar distance remains sufficiently short after the second CT, the emerging N+ ion can bind to the neutral Ar atom, yielding a new [NAr]+ species via ion transfer. This secondary surface crossing therefore opens two additional dissociation pathways:
N2Ar2+ NAr+ NAr N+ + N+ + Ar III
N2Ar2+ NAr+ NAr N+ + [NAr]+ IV
@@
As we show below, these fragmentation channels occur exclusively following photoionization of the Ar 2p shell. Notably, while Fig. 1 illustrates crossings among three representative electronic states, ab initio calculations of the adiabatic potential energy curves reveal dozens of intersecting states in this energy region, providing multiple pathways for both SCT and DCT processes.
3 Experimental Results and Discussion
@@
In Fig. 2 we present the time-of-flight correlation of two ions measured in coincidence with an electron. Panel (a) shows our results for a photon energy of eV, while we used photons with an energy of 420 eV for the initial ionization of our gas target in Panel (b). In addition, we filtered both panels by requiring a photoelectron originating from the Ar and N2 sites, respectively. A Coulomb explosion involving a breakup into exactly two ionic fragments – such as N + Ar+, N+ + [NAr]+ (and N+ + N+ resulting from the ionization of uncondensed N2 molecules) – the fragments are emitted back-to-back (i.e., with momenta of equal magnitude but opposite direction) due to momentum conservation, resulting in a sum momentum of the fragments of zero. This yields narrow diagonal features in the coincidence maps shown in Fig. 2 and used to identify the ion pairs occurring in the final state. In contrast, the N+ + Ar+ + N channel appears as a broader structure, since the undetected neutral N atom carries away a part of that sum momentum. In addition, the dissociation channel N+ + N+ + Ar is expected to overlap with the aforementioned N+ + N+-channel, but—as we show below—can be disentangled by our coincidence measurement.
@@
We observe dimer fragmentation features belonging to all processes listed in Section 3.2 and labeled them accordingly in Figs. 2(a) and 2(b). In particular, we observe the breakup of the dimer after Ar 2p-ionization into N+ + [NAr]+, i.e., events belonging to channel IV that requires a double charge transfer and an ion transfer.
3.1 Single Charge Transfer Pathways
To elucidate the mechanisms by which an electron is transferred across the dimer, we analyze the kinetic energy release (KER) spectra of the ionic fragments. Figure 3(a) shows the measured KER distribution for the N + Ar+ channel associated with the SCT process, featuring a dominant peak at about 4.7 eV and a weaker peak at around 3.5 eV. In weakly bound systems, the KER resulting from a Coulomb explosion can be approximated by the Coulomb potential of two point-like charges that are located at a given internuclear distance . Accordingly, within this approximation it is possible to extract internuclear distances from the measured KER. A KER of 3.5 eV corresponds to R = , consistent with the equilibrium distance of R = [37].
@@ 3.1 Single Charge Transfer Pathways
By contrast, the 4.7 eV peak corresponds to R = , where N2Ar2+ ions produced via Auger decay occupy an attractive region of the PES and SCT occurs at an internuclear distance significantly shorter (by ) than equilibrium [38]. This finding indicates that the CT proceeds on a timescale comparable to, or slower than, the vibrational motion of the cluster, consistent with the relatively large separation between constituents. Moreover, the N fragment in the N + Ar+ channel can reside in a dissociative state, yielding the N+ + Ar+ + N channel. Figure 3(b) presents the corresponding KER distribution of the N+ and Ar+ fragments, which exhibits two peaks at 4.5 eV and 2.4 eV. The higher-energy peak reflects direct fragmentation of N following SCT, where essentially the entire Coulomb potential is shared between N+ and Ar+, while the neutral N atom carries negligible kinetic energy. The resulting total KER closely matches the main peak in Fig. 3(a). In contrast, the lower-energy peak arises when N is at first repelled by the Ar+ as an intact entity and subsequently dissociates, with the neutral N atom removing about half the energy initially carried by N. This redistribution produces a smaller KER with a broader width.
@@
To test whether the CT mechanism persists in more complex scenarios, we selectively ionized the N2 site in N2Ar dimers using photons of 420 eV. Following N 1s photoionization, Auger decay produces doubly charged NAr ions. Figure 3(c) shows the measured KER distribution for the N + Ar+ channel resulting from N2 site ionization. A KER of 3.5 eV corresponds to R = , again consistent with the equilibrium separation of R = , as illustrated in Fig. 1. Likewise, N ions formed from NAr via SCT can further dissociate into N+ + N fragments, corresponding to pathway II in Fig. 1. As shown in Fig. 3(d), the KER distribution of N+ and Ar+ exhibits lower values and broader widths compared with the N + Ar+ channel. This broadening arises because N may begin to dissociate and release energy prior to the SCT event.
3.2 Double Charge Transfer Pathways
@@
4 Conclusion
In conclusion, we investigated ultrafast multiple CT processes mediated by nonadiabatic electron-nuclear dynamics in weakly bound systems using state-of-the-art multi-coincidence spectroscopy in combination with ab initio calculations. Our investigations were performed using a prototype system that combines aspects of van der Waals and covalent bonding: The N2Ar dimer is complex enough to enable even double charge transfer processes and intriguing ultrafast molecular dynamics. Yet, its complexity is still suitable for a detailed study of these processes on a single-atom level. The intermolecular relaxation was initiated by site-selective ionization of the N2 molecule or the Ar atom within the dimer using photons in the X-ray and XUV regime. Both SCT and DCT processes were unambiguously identified experimentally and theoretically starting from well-defined initial states, namely N2Ar2+ and N22+Ar. The DCT pathways were experimentally established by observing the N+ + N+ + Ar and N+ + [NAr]+ channels in coincident electron and ion energy spectra. For the latter case, a heavy N+ ion transfer channel was identified leading to the formation of [NAr]+. This channel, observed experimentally in combination with calculations, is induced by DCT processes through two conical intersections in the N2Ar2+ initial state but not in NAr. This mechanism fundamentally alters the character of CT in comparison to that of atom-atom dimers or clusters, where CT typically occurs only once. Our results thus demonstrate that sequential multiple CT processes, coupled to electron-nuclear dynamics, can drive ion transfer even in the simplest molecular aggregates. DCT was not observed in doubly charged NAr dimers, highlighting that rapid dissociation of molecular ions can suppress CT dynamics. Altogether, this study provides new insights into nonadiabatic intermolecular (multiple) CT dynamics, showing that even a simple single-molecule mixed-interaction system can undergo complex reaction pathways – an observation of broad relevance for understanding functional processes in complex biochemical systems.
5 Methods
@@
The experiments were performed using two synchrotron light sources at different photon energies: 260 eV and 420 eV were employed at the soft X-ray beamline P04 of the PETRA III synchrotron (DESY, Hamburg, Germany) in timing mode (40 bunches, 192 ns bunch separation) with circularly polarized light [41]; meanwhile, a photon energy of 43.1 eV was used at the Flexible PhotoElectron Spectroscopy (FlexPES) beamline of MAX IV (Lund, Sweden) in single-bunch mode (320 ns bunch separation) with linearly polarized light [42]. In both facilities, a supersonic gas jet was formed, collimated by two skimmers, and crossed with the photon beam inside a COLTRIMS reaction microscope [33, 43, 34], with N2Ar dimers produced by expanding a 50/50 gas mixture through a 30 m nozzle at 11 bar (PETRA III) or a 50 m nozzle at 5 bar (MAX IV), both at 300 K. The employed COLTRIMS spectrometers consisted of an ion arm of 31 mm (PETRA III) or 34 mm (MAX IV) acceleration length and an electron arm of 70 mm acceleration and 153 mm drift length (Wiley-McLaren time-of-flight focusing geometry, PETRA III) or 419 mm acceleration length (MAX IV). All were equipped with microchannel plate detectors (active area of 120 mm diameter for the electron detector at PETRA III and of 80 mm diameter for all other detectors) with hexagonal delay-line position readout [44, 45]. Electrons and ions were guided to position-sensitive microchannel plate detectors with delay-line readout using optimized electric and magnetic fields, 20.8 V/cm and 5.6 G at PETRA III, and 12.4 V/cm and 7.7 G at MAX IV, allowing for the complete 4 collection of electrons with kinetic energies up to 30 eV and 23 eV, respectively. Data were recorded in list mode files, measuring the positions and times of flight of all reaction products for each ionization event; this enabled the extraction of the weak dimer signals and the reconstruction of particle momenta in offline analysis, despite most events originating from monomer ionization, and charge transfer channels were ultimately identified by selecting events with two ions of equal and opposite momentum, whose sharp back-to-back emission line in the PIPICO spectrum provides a unique signature of the final Coulomb explosion step.
@@
Acknowledgements This work was supported by the National Key Research and Development Program of China (No. 2022YFA1604302), the National Natural Science Foundation of China (No. 12334011, No. 12204498, No. 12474259, No. 12374262, and No. 12450404), and the Natural Science Foundation of Henan (No. 252300421304). F.T. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project 509471550, Emmy Noether Programme.
@@
Conflict of interest/Competing interests (check journal-specific guidelines for which heading to use)

### v2

@@ Abstract
Ultrafast charge transfer (CT) processes redistribute electronic charge within and between molecular units and play a central role in many physical, chemical, and biological phenomena. However, the microscopic pathways of multiple CT events, including the coupled structural evolution and energy redistribution, are challenging to disentangle experimentally in complex systems. To obtain controlled insight into such dynamics, well-defined properties are required. Here, we investigate the N2Ar dimer, which combines a covalent bond with a weak van der Waals interaction, using site-selective synchrotron photoionization and coincident detection of electrons and ions. Combined with ab initio calculations, this approach enables step-by-step tracking of ultrafast CT and fragmentation dynamics. We find that the dimer’s structural evolution triggers a second CT event, opening complex reaction pathways in which electrons are transferred back and forth between Ar and N2, through two nonadiabatic transitions involving conical intersections. These results demonstrate that sequential multiple CT-induced transitions, even in a simple dimer, provide controlled insight into nonadiabatic reaction mechanisms relevant to complex systems.
@@
Ultrafast charge transfer (CT), which redistributes electronic charge within or between molecular units, is a fundamental process underlying functionality in a wide range of natural and engineered systems [1, 2, 3, 4, 5, 6, 7]. Charge and energy transport span length scales from small dimers and isolated molecules in the gas or liquid phase to photoactive materials and proteins [8, 9, 10, 11, 12, 13]. Tracing complex reactions induced by multiple CT via bulk ensemble measurements is often challenging, and key microscopic aspects of these processes can remain difficult to resolve. Access to well-defined molecular systems therefore provides an important route to uncover transient stages of CT dynamics and to elucidate the elementary quantum motion of electrons and nuclei involved in these processes [14, 15].
@@
In practice, charge and energy exchange typically occur in complex environments where covalent interactions coexist with weaker noncovalent forces such as van der Waals bonding. The latter, despite being the weakest intermolecular interaction, plays a crucial role in determining the structure and dynamics of molecular aggregates, condensed phases, and biological matter [16, 17, 19, 20, 18]. Simple rare-gas dimers, however, are not ideal systems for studying complex multiple CT-driven reaction dynamics, as they usually dissociate promptly after a single CT event [21, 22]. A suitable prototype system is the N2Ar dimer, which features a strong covalent bond within N2 and a weak van der Waals bond between N2 and Ar. Its distinct properties make it a prototypical emulating model for experiment and quantum-chemical calculations to probe reaction mechanisms relevant to complex systems. First, the nearly matching ionization potentials [23, 24] of its two constituents make intermolecular reactions more effective under photoionization and photoexcitation scenarios by creating numerous crossings on the potential energy surfaces. Second, its structural flexibility, especially the stretching of its bonds [25] enables access to a manifold of nonadiabatic pathways through conical intersections and initiates the occurrence of multiple CT. Therefore, a system as simple as N2Ar can effectively mimic the ensemble reactions occurring in complex condensed chemical and biological systems.
@@
X-ray-induced inner- or core-shell ionization offers an element-specific route to prepare this system in well-defined charge states. In this work, we investigate experimentally and theoretically the photoionization-induced nonadiabatic dynamics in N2Ar, focusing on ultrafast multiple CT processes. Using datasets recorded at different photon energies, we selectively initiate double ionization (DI) of either N2 or Ar via Auger decay or direct double ionization, thereby preparing well-defined initial states. By analyzing the momenta of dimer fragments and electrons in the final state, we have unambiguously disentangled various CT pathways, enabling reconstruction of fragmentation pathways driven by SCT and double CT (DCT). Our findings demonstrate that even in such a simple dimer multiple CT-induced transitions can drive complex reaction channels, providing fundamental insight into charge redistribution as a core nonadiabatic mechanism underlying functional processes in complex natural systems.
@@
Expected Charge Transfer Channels and Involved Potential Energy Surfaces
@@
To understand how the observed reaction channels and ultrafast CTs evolve from the doubly charged N2Ar dimer, we calculated the corresponding potential energy surfaces (PESs) and traced possible fragmentation pathways. Details of the calculations are in the Methods section. For [N2Ar]2+, charge localization can occur in three configurations: NAr, N2Ar2+, and NAr+. The high density of electronic states and the vibrational/rotational degrees of freedom of N2 lead to significant overlap among the PESs over a wide range of N2Ar bond lengths and angles. This extensive overlap facilitates the exceptionally rapid nonadiabatic CT dynamics, which is rare in simpler systems like Ar2 dimers. Selective ionization using X-ray photons prepares specific initial states: Auger decay following removal of an N 1s or Ar 2p electron generates NAr (e.g., , a, ) or N2Ar2+ (e.g., , , ) states, respectively [31, 32]. Intermolecular interactions then drive the cluster ion’s structural evolution, potentially leading to a nearly linear N2Ar geometry.
@@
Figure 1 illustrates the two-dimensional PESs for three electronic states of the linear [N2Ar]2+ ion: N(a)Ar, NAr2+(), and N(D)Ar+(). When the system starts in N(a)Ar or N2Ar2+(), the N-N vibrational wave packet lies within the crossing region with the N(D)Ar+() surface, enabling prompt SCT to form the N + Ar+ ion pair. The N-N bond elongates because the equilibrium bond length of N(D) is much larger than that of neutral N2 or N(a). Coulomb repulsion then stretches the N2-Ar bond, promoting dissociation along two distinct SCT pathways:
N2Ar2+ / NAr NAr+ N + Ar+ (I)
N2Ar2+ / NAr NAr+ N+ + N + Ar+ (II)
@@
As the N-N bond continues to elongate, the N(D)Ar+() and N(a)Ar PESs intersect again, allowing a second CT event. For clusters starting in N(a)Ar, dissociation is two-step: N(a) captures an electron from Ar, forming NAr+, and then, as the N-N bond stretches, an electron transfers back to Ar+. If the initial state is N2Ar2+(), neutral N2 sequentially accepts two electrons from Ar2+, ultimately driving dissociation. Crucially, if the N2-Ar distance is sufficiently short after the second CT, the emerging N+ ion can bind to neutral Ar, yielding the new [NAr]+ species via ion transfer. This secondary surface crossing opens two additional dissociation pathways:
N2Ar2+ NAr+ NAr N+ + N+ + Ar (III)
N2Ar2+ NAr+ NAr N+ + [NAr]+ (IV)
@@
As we show below, these fragmentation channels (III and IV) occur exclusively following photoionization of the Ar 2p shell. While Fig. 1 shows crossings among three representative states, ab initio calculations reveal dozens of intersecting states, providing multiple SCT and DCT pathways.
Experimental Results and Discussion
@@
In Fig. 2 we present the time-of-flight correlation of two ions measured in coincidence with an electron. Panel (a) shows our results for a photon energy of eV, while we used photons with an energy of 420 eV for the initial ionization of our gas target in Panel (b). In addition, we filtered both panels by requiring a photoelectron originating from the Ar and N2 sites, respectively. A Coulomb explosion involving a breakup into exactly two ionic fragments – such as N + Ar+, N+ + [NAr]+ (and N+ + N+ resulting from the ionization of uncondensed N2 molecules) – the fragments are emitted back-to-back (i.e., with momenta of equal magnitude but opposite direction) due to momentum conservation, resulting in a sum momentum of the fragments of zero. This yields narrow diagonal features in the coincidence maps shown in Fig. 2 and used to identify the ion pairs occurring in the final state. In contrast, the N+ + Ar+ + N channel appears as a broader structure, since the undetected neutral N atom carries away a part of that sum momentum. In addition, the dissociation channel N+ + N+ + Ar is expected to overlap with the aforementioned N+ + N+ channel, but—as we show below—can be disentangled by our coincidence measurement.
@@
We observe a dimer fragmentation features belonging to all processes listed in Section Double Charge Transfer Pathways and labeled them accordingly in Figs. 2(a) and 2(b). In particular, we observe the breakup of the dimer after Ar 2p ionization into N+ + [NAr]+, i.e., events belonging to channel IV that requires a double charge transfer and an ion transfer.
Single Charge Transfer Pathways
@@
To elucidate the SCT mechanisms, we analyzed the kinetic energy release (KER) spectra of the ionic fragments. For the N + Ar+ channel [Ar site ionization, Fig. 3(a)], the KER distribution features a dominant peak at 4.7 eV and a weaker peak at 3.5 eV. Using the point-charge Coulomb approximation, a KER of 3.5 eV corresponds to an internuclear distance R = , consistent with the equilibrium distance of R = [37]. The 4.7 eV peak corresponds to R = , indicating that N2Ar2+ ions are produced in an attractive region of the PES, and SCT occurs at a significantly shorter distance [38]. This suggests the CT timescale is comparable to, or slower than, the cluster’s vibrational motion. The N fragment can also be in a dissociative state, yielding the N+ + Ar+ + N channel. The corresponding KER distribution [Fig. 3(b)] exhibits peaks at 4.5 eV and 2.4 eV. The higher-energy peak reflects direct N fragmentation following SCT, where the total KER closely matches the main N + Ar+ peak. The lower-energy, broader peak arises when N is repelled by Ar+ before dissociation, and the neutral N atom carries away part of the energy.
@@
To test whether the CT mechanism persists in more complex scenarios, we selectively ionized the N2 site in N2Ar dimers using photons of 420 eV. Following N 1s photoionization, Auger decay produces doubly charged NAr ions. The measured KER for the N + Ar+ channel [Fig. 3(c)] shows a peak at 3.5 eV, consistent with the equilibrium separation. The N+ + Ar+ KER distribution [Fig. 3(d)] shows lower values and broader widths compared with the N + Ar+ channel. This broadening occurs because N may begin to dissociate and release energy before the SCT event.
Double Charge Transfer Pathways
@@
In conclusion, we investigated ultrafast multiple charge transfer (CT) processes mediated by nonadiabatic electron-nuclear dynamics in a weakly bound molecular system using multi-coincidence spectroscopy in combination with ab initio calculations. The N2Ar dimer, which combines covalent and van der Waals bonding, serves as a well-defined prototype that is sufficiently complex to support sequential CT processes while remaining amenable to detailed experimental and theoretical analysis. Site-selective ionization of either the N2 molecule or the Ar atom initiates intermolecular relaxation dynamics from well-defined initial charge states. Both single and double CT pathways were identified experimentally and supported by theoretical modeling, starting from the dicationic states N2Ar2+ and N22+Ar. Double CT processes were unambiguously established by observing the N+ + N+ + Ar and N+ + [NAr]+ fragmentation channels in coincident electron-ion measurements. In particular, the formation of [NAr]+ was shown to result from a sequential double CT mechanism involving two nonadiabatic transitions at conical intersections, rather than from direct dissociation or heavy-ion tunneling. This mechanism leads to an effective ion-transfer process that qualitatively differs from the single CT dynamics typically observed in simpler atom-atom dimers. Our results demonstrate that sequential multiple CT events, coupled to ultrafast structural rearrangement, can drive complex reaction pathways even in a minimal mixed-interaction system. At the same time, the absence of double CT following ionization at the N2 site highlights how rapid molecular dissociation can suppress CT dynamics. More generally, these findings illustrate how controlled model systems can be used to disentangle nonadiabatic charge redistribution mechanisms and to provide insight into CT-driven reaction dynamics that are relevant to a broad range of molecular and condensed-phase environments.
Methods
@@
The experiments were performed using two synchrotron light sources at different photon energies: 260 eV and 420 eV were employed at the soft X-ray beamline P04 of the PETRA III synchrotron (DESY, Hamburg, Germany) in timing mode (40 bunches, 192 ns bunch separation) with circularly polarized light [41]; meanwhile, a photon energy of 43.1 eV was used at the Flexible PhotoElectron Spectroscopy (FlexPES) beamline of MAX IV (Lund, Sweden) in single-bunch mode (320 ns bunch separation) with linearly polarized light [42]. In both facilities, a supersonic gas jet was formed, collimated by two skimmers, and crossed with the photon beam inside a COLTRIMS reaction microscope [33, 43, 34], with N2Ar dimers produced by expanding a 50/50 gas mixture through a 30 m nozzle at 11 bar (PIPE endstation PETRA III) or a 50 m nozzle at 5 bar (ICE endstation MAX IV), both at 300 K. The employed COLTRIMS spectrometers consisted of an ion arm of 31 mm (PIPE endstation) or 34 mm (ICE endstation) acceleration length and an electron arm of 70 mm acceleration and 153 mm drift length (Wiley-McLaren time-of-flight focusing geometry, PIPE endstation) or 419 mm acceleration length (ICE endstation). All were equipped with microchannel plate detectors (active area of 120 mm diameter for the electron detector at PETRA III and of 80 mm diameter for all other detectors) with hexagonal delay-line position readout [44, 45]. Electrons and ions were guided to position-sensitive microchannel plate detectors with delay-line readout using optimized electric and magnetic fields, 20.8 V/cm and 5.6 G at PETRA III, and 12.4 V/cm and 7.7 G at MAX IV, allowing for the complete 4 collection of electrons with kinetic energies up to 30 eV and 23 eV, respectively. Data were recorded in list mode files, measuring the positions and times of flight of all reaction products for each ionization event; this enabled the extraction of the weak dimer signals and the reconstruction of particle momenta in offline analysis, despite most events originating from monomer ionization, and charge transfer channels were ultimately identified by selecting events with two ions of equal and opposite momentum, whose sharp back-to-back emission line in the PIPICO spectrum provides a unique signature of the final Coulomb explosion step.
@@
Acknowledgements This work was supported by the National Key Research and Development Program of China (No. 2022YFA1604302), the National Natural Science Foundation of China (No. 12334011, No. 12541505, No. 12204498, No. 12474259, No. 12374262, and No. 12450404), and the Natural Science Foundation of Henan (No. 252300421304). F.T. acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project 509471550, Emmy Noether Programme.
@@
Conflict of interest/Competing interests: